OpenStudy (anonymous):

Simplify one fourth(g − 16)

OpenStudy (hba):

\[\frac{ 1 }{ 4 }(g-16)\]

OpenStudy (hba):

\[\frac{ g }{ 4 }-4\]

OpenStudy (anonymous):

way to explain it

OpenStudy (anonymous):

that's not even an answer choice.. :(

OpenStudy (amorfide):

there is no explanation really, just multiply out lol

OpenStudy (hba):

\[\frac{ g-16 }{ 4}\]

OpenStudy (amorfide):

@rachelraver what are the annswers?

OpenStudy (anonymous):

ummm, brittany it's not 1/4 = g - 16

OpenStudy (anonymous):

1/4g − 16 1/4g − 4 1/4g + 4 1/4g + 16

OpenStudy (hba):

\[\frac{ 1 }{ 4 }g-4\]

OpenStudy (hba):

Which Is Just The Same As I Showed Up LoL

OpenStudy (hba):

In My Second Comment

OpenStudy (amorfide):